Nonconstant Growth Valuation

A company currently pays a dividend of $1.75 per share (D0 = $1.75). It is estimated that the company's dividend will grow at a rate of 18% per year for the next 2 years, and then at a constant rate of 5% thereafter. The company's stock has a beta of 2, the risk-free rate is 3%, and the market risk premium is 3%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Solution

For estimating the stock's current price, we have to compute the Cost of Capital.

Cost of capital (under CAPM Model) = Risk-free rate + (Beta X Risk Premium)

Here, Risk-free rate = 3%

Beta = 2

Risk Premium = 3%

Therefore, Cost of Capital (CAPM) = 3% + (2 X 3%) = 9%

Now, let us value stock's current price,

Dividend paid (D0) = $ 1.75 per share

Dividend will grow @ 18%, or 0.18 per $ for next 2 years, and then constant rate of 5%, or 0.05 per $ thereafter

Therefore, D1 = D0 X ( 1 + g) = $ [ 1.75 X (1 + 0.18) ] = $ 2.065

D2 = D1 X ( 1 + g) = $ [ 2.065 X (1 + 0.18) = $ 2.4367

D3 = D2 X (1+g) = $ [ 2.4367 X (1+0.05)] = $ 2.558535

Now, P2 = D3 / (Cost of Capital - g) = $ [ 2.558535 / (0.09 - 0.05)] = $ 63.96 (Approx)

Now, Current Market Price (P0) = [D1/(1+Cost of Capital)] + [(D2+P2)/(1+Cost of Capital)2]

=$ [2.065/(1+0.09)] + [(2.4367+63.96)/(1+0.09)2]

= $ (1.89 + 55.89)

= $ 57.78

Answer: Expected Current Market Price of the stock is $ 57.78.